OpenStudy (anonymous):

Find all the solutions to: xdot = [1 2; 0 1] * x

myininaya (myininaya):

can you tell me what this means

OpenStudy (anonymous):

Sure! I'm asking how to solve a system of linear differential equations of the form: \[x \prime = \left[\begin{matrix}1 & 2 \\ 0 & 1\end{matrix}\right] * \left(\begin{matrix}x1 \\ x2\end{matrix}\right)\]

myininaya (myininaya):

oh thats x prime?

myininaya (myininaya):

do you know the steps?

OpenStudy (anonymous):

So from what I understand, I first have to find the eigenvalues of the matrix \[\left[\begin{matrix}1 & 2 \\ 0 & 1\end{matrix}\right]\] , which gives me det(\[\left[\begin{matrix}1-\lambda & 2 \\ 0 & 1-\lambda\end{matrix}\right]\]) = 0. Solving this gives me \[\lambda = 1\], with a multiplicity of two. I'm confused as to how to capture the effect of multiplicity.

myininaya (myininaya):

i'm looking for an example

OpenStudy (anonymous):

Do you have a Cramster account? I've found a good example there if you can't find one. The problem is, it's not helping me as much as I would've hoped.

myininaya (myininaya):

ok it looks like we also need to find eigenvectors

OpenStudy (anonymous):

Right, and when I solve for the eigenvectors I get v1 = \[\left(\begin{matrix}0 \\ 1\end{matrix}\right)\] And because there is a double root, it's the only eigenvector found. From what I've read, the multiplicity adds other terms but that's where my confusion lies.

myininaya (myininaya):

why is it so hard to find an example lol

OpenStudy (anonymous):

Haha, try this one! http://www.cramster.com/solution/solution/1085373

myininaya (myininaya):

okay i think this is good i'm gonna try to work your problem

OpenStudy (anonymous):

Thank you! I'd be grateful if you could get me out of this rut.

myininaya (myininaya):

hmmm... they were about to find another vector that wasn't linear dependent

myininaya (myininaya):

i don't see how we can find a vector that is linear independent from the first one

OpenStudy (anonymous):

I think the multiplication by 't' makes it linearly independent, right?

myininaya (myininaya):

thats really the only thing thats holding us back

OpenStudy (anonymous):

Would you happen to know anyone else that help us? I'm not sure this should be this difficult.

OpenStudy (phi):

the eigenvectors are [1 0] and [0 1]

OpenStudy (anonymous):

Hi phi! Would you mind telling me how you got that?

myininaya (myininaya):

i don't know how he got that but i do have something but i don't know if it is right

myininaya (myininaya):

i will rewrite it

OpenStudy (anonymous):

Alright!

myininaya (myininaya):

i got these two vectors for lambda=1 [1 0] and [1 1/2] i will show you how in a second

OpenStudy (anonymous):

Alright that would be great, I'm not sure how you got those.

OpenStudy (phi):

oops I lied, the evecs are [1 0] and [-1 0]

myininaya (myininaya):

but aren't those linear dependent phi?

OpenStudy (anonymous):

Hmm, alright I think I get it. I'm just going to go through your work once more really quickly. Thank you again for taking the time to answer it!

myininaya (myininaya):

well like i said i don't know if it is right

OpenStudy (anonymous):

so I guess I'm confused by how the first one you deduced that p1 could be anything?

OpenStudy (anonymous):

Oh because it is multiplied by zero?

myininaya (myininaya):

yep

OpenStudy (phi):

If you guys find a second eigenvector please let me know!

OpenStudy (anonymous):

Yeah I understand this method, I hope it's right! I'm going to think about it some more.

OpenStudy (zarkon):

I get \[c_1\left[\begin{matrix}1 \\ 0\end{matrix}\right]e^t+c_2\left(\left[\begin{matrix}1 \\ 0\end{matrix}\right]t+\left[\begin{matrix}0 \\ 1/2\end{matrix}\right]\right)e^t\]

OpenStudy (anonymous):

Hi Zarkon, could you possibly explain how you got that?

OpenStudy (zarkon):

I just followed the example